1 1/15 + 3 3/10 - 2 4/5 = ??
Can someone help me with this please, I really have no idea where to start.
Thank you.

Answer by edjones(8007)

1 1/15 + 3 3/10 - 2 4/5
=16/15+33/10-14/5
=32/30+99/30-84/30
=47/30 or 1 17/30
Ed